Probability

Get insights from 304 questions on Probability, answered by students, alumni, and experts. You may also ask and answer any question you like about Probability

  • 304

    Questions

  • 0

    Discussions

  • 0

    Active Users

  • 4

    Followers

Have a question about Probability? Ask our experts

Ask Your Question

Keep it short & simple. Type complete word. Avoid abusive language.

New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

Let      E1:The  event  that  the  letter  comes  from  TATANAGARand     E2:The  event  that  the  letter  comes  from  CALCUTTAAlso,  E3:The  event  that  on  the  letter,  two    letters  TA  are  visible.∴P(E1)=12,  P(E2)=12  and  P(E3E1)=28  and  P(E3E2)=17[?  For  TATA  NAGAR,  the  two consecutive   letters  visible  are  TA,AT,TA,AN,NA,AG,GA,AR]∴  P(E3/E1)=28and  [For  CALCUTTA,  the  two consecutive   letters  visible  are  CA,AL,LC,CU,UT,TT  and  TA]So,  P(E3/E2)=17Now  using  Bayes'  Therorm,  we  have∴P(E1/E3)=P(E1).P(E3/E1)P(E1).P(E3/E1)+P(E2).P(E3/E2)                  =12.2812.28+12.17=1818+114=187+456=711Hence,  the  required  probability  is  711.

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

Given  that  A1:A2:A3=4:4:2∴  P(A1)=410,  P(A2)=410  and  P(A3)=210  whereA1,A2  and  A3  are  the  three  types  of  seeds.Let  E  be  the  events  that  a  seed    and  E¯  be  the  events  that  a  seed  does  not  ∴  P(EA1)=45100,  P(EA2)=60100  and  P(EA3)=35100and  P(E¯A1)=55100,  P(E¯A2)=40100  and  P(E¯A3)=65100(i)  P(E)=P(A1).P(EA1)+P(A2).P(EA2)+P(A3).P(EA3)                     =410.45100+410.60100+210.35100                     =1801000+2401000+701000=4901000=0.49(ii)  P(E¯/A3)=1−P(E/A3)=1−35100=65100=0.65(iii)  ,  Bayes'  Theorem,  we  get         P(A2/E¯)=P(A2).P(E¯/A2)P(A1).P(E¯/A1)+P(A2).P(E¯/A2)+P(A3).P(E¯/A3)                              =410.40100410.55100+410.40100+210.65100=16010002201000+1601000+1301000                              =160220+160+130=160510=1651=0.314Hence,  the  required  probability  is  1651  or  0.314

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

  Referring to Exercise  Q.41,  we  will  use  here,  Bayes'  Theorem(i)  P(E2/F)=P(E2).P(F/E2)P(E1).P(F/E1)+P(E2).P(F/E2)+P(E3).P(F/E3)                             =26.1316.0+26.13+36.1=218218+36=218×1811=211(ii)  P(E3/F)=P(E3).P(F/E3)P(E1).P(F/E1)+P(E2).P(F/E2)+P(E3).P(F/E3)                              =36.116.0+26.13+36.1=36218+36=36×1811=911Hence,  the  required  probabilities  are  211  and  911

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Long Answer Type Questions as classified in NCERT Exemplar

Sol:

Given  that:       ???   Bag I   =3  red  balls  and  no  white  ball          Bag II=2  red  balls  and  1  white  ball         Bag III=no  red  ball  and  3  white  ballsLet  E1,  E2  and  E3  be  the  events  of    Bag II,  and  Bag III  respectively  and  a  ball  isdrawn  from  it.∴P(E1)=16,  P(E2)=26  and  P(E3)=36(i)  Let  E  be  the  event  that  red  ball  is  selected∴P(E)=P(E1).P(E/E1)+P(E2).P(E/E2)+P(E3).P(E/E3)                  =16.33+26.23+36.0=318+418=718(ii)  Let  F  be  the  event  that  white  ball  is  selected∴P(F)=1−P(E)                                [P(E)+P(F)=1]                  =1−718=1118Hence,  the  required  probabilities  are  718  and  1118.

Thumbs Up IconUpvote Thumbs Down Icon
New question posted•a year ago
New answer posted•a year ago
S
a year ago

95.

P (A)+P (B)−P (AandB)=P (A)⇒P (A)+P (B)−P (A∩B)=P (A)⇒P (B)−P (A∩B)=0⇒P (A∩B)=P (B)∴P (A|B)=P (A∩B)P (B)=P (B)P (B)=1

Therefore, option (B) is correct.

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

94.

P (A|B)>P (A)⇒P (A∩B)P (B)>P (A)⇒P (A∩B)>P (A).P (B)⇒P (A∩B)P (A)>P (B)⇒P (B|A)>P (B)

Therefore, option (C) is correct.

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

93. P (A)≠0 and P (B|A)=1

P (B∩A)=P (B∩A)P (A)1=P (B∩A)P (A)P (A)=P (B∩A)⇒A⊂B

Therefore, option (A) is correct.

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

92. Let E1 = Ball transferred from Bag I to Bag II is red

E2 = Ball transferred from Bag I to Bag Ii is black

A = Ball drawn from Bag II is red in colour

P (E1) =3/7 and P (E2) = 4/7

Let A be the event that the ball drawn is red.

When a red ball is transferred from bag I to II,

P (A | E1) = 5/10 = 1/2

Whe

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

91. Let the event in which A fails and B fails be denoted by EA and EB.

P (EA) = 0.2

P (EA ∩ EB) = 0.15

P (B fails alone) = P (EB) − P (EA ∩ EB)

⇒ 0.15 = P (EB) − 0.15

⇒ P (EB) = 0.3

P (EA|EB)=P (EA∩EB)P (EB)=0.150.3=0.5

(ii) P (A fails alone) = P (EA) − P (EA ∩ EB)

= 0.2 − 0.15

= 0.05

Thumbs Up IconUpvote Thumbs Down Icon